# DialSim: A Real-Time Simulator for Evaluating Long-Term Multi-Party Dialogue Understanding of Conversation Systems

### v5

@@ Abstract
Recent advancements in Large Language Models (LLMs) have significantly enhanced the capabilities of conversation systems, making them applicable to various fields (e.g., education). Despite their progress, the evaluation of the systems often overlooks the complexities of real-world conversations, such as real-time interactions, multi-party dialogues, and extended contextual dependencies. To bridge this gap, we introduce DialSim, a real-time dialogue simulator. In this simulator, a conversation system is assigned the role of a character from popular TV shows, requiring it to respond to spontaneous questions using past dialogue information and to distinguish between known and unknown information. Key features of DialSim include assessing the system’s ability to respond within a reasonable time limit, handling long-term multi-party dialogues, and evaluating performance under randomized questioning with LongDialQA, a novel, high-quality question-answering dataset. Our experiments using DialSim reveal the strengths and weaknesses of the latest conversation systems, offering valuable insights for future advancements in conversational AI. DialSim is available at https://dialsim.github.io/.
DialSim: A Real-Time Simulator for Evaluating Long-Term
Multi-Party Dialogue Understanding of Conversation Systems
Jiho Kim1, Woosog Chay1, Hyeonji Hwang1, Daeun Kyung1, Hyunseung Chung1, Eunbyeol Cho1, Yohan Jo2, Edward Choi1 1KAIST  2SNU {jiho.kim, edwardchoi}@kaist.ac.kr
@@ 1 Introduction
Figure 1: An overall process of DialSim. Gray speech bubbles indicate predetermined utterances from the script, and white speech bubbles indicate spontaneous questions asked during the simulation. Colored speech bubbles indicate the system’s responses to the questions. (Left) An unanswerable question. (Center) A question that references a specific time. (Right) A multi-hop question that requires understanding past sessions (i.e., the Left and Center boxes). The question is asked in the format chosen by the user, either in a multiple-choice format or as an open-ended question.
@@
Recent advancements in Natural Language Generation (NLG) within Large Language Models (LLMs) have significantly enhanced the capabilities of conversation systems. These systems are now integral to various fields, including entertainment (Zhou et al., 2023; Chen et al., 2024) and education (Ait Baha et al., 2023; Waisberg et al., 2024), providing personalized interactions that cater to individual preferences and interests. As they continue to evolve and become more widely adopted, it is crucial to rigorously assess their performance in real-world scenarios to ensure they meet user expectations and function effectively.
Traditionally, the evaluation of conversation systems has relied on qualitative assessments of their responses. This process typically involves human evaluators or LLMs judging the quality of the system’s responses (Adiwardana et al., 2020; Zhang et al., 2020; Roller et al., 2021; Shuster et al., 2022; Lee et al., 2023; Kim et al., 2024) or comparing responses between different systems on platforms like Chatbot Arena (Chiang et al., 2024). While these methods provide valuable insights into aspects such as naturalness and alignment with user instructions, they do not fully capture the complexities of real-world interactions.
In practical scenarios, conversation systems must handle real-time interactions, manage multi-party conversations, and recall relevant information from past dialogues. These challenges demand more comprehensive evaluation methods—ones that test a system’s ability to respond within a reasonable time constraint, understand multi-party dialogue contexts, and reason across extended interactions. To address this need, we introduce DialSim, a real-time dialogue simulator designed to evaluate the long-term multi-party dialogue understanding of conversation systems.
DialSim places the conversation system in the role of a main character within a TV show, engaging in extensive conversations based on the show’s scripted content (see Figure 1). During each session, a randomly selected character asks a randomly sampled question at an unpredictable time. The system is evaluated on its ability to respond appropriately, relying solely on the dialogue history and acknowledging when it lacks sufficient information. This approach enables rigorous testing of dialogue comprehension in unpredictable, realistic scenarios. Additionally, the system’s real-time interaction capabilities are assessed through time constraints for responses (e.g., 3s, 5s). Note that DialSim is designed not to evaluate a single LLM, but to evaluate a conversation system’s dialogue understanding ability in an environment as realistic as possible.
@@
To implement DialSim, a dialogue script and corresponding question-answer (QA) pairs are required. For this purpose, we created LongDialQA, a new QA dataset derived from long-term multi-party dialogues. It comprises dialogues from popular TV shows (i.e., Friends, The Big Bang Theory, and The Office), covering 1,300 sessions over five years, totaling around 350,000 tokens. Each session includes more than 1,000 questions curated through two approaches: refining questions from a fan quiz website and generating complex questions using extracted temporal knowledge graphs.
To rigorously evaluate the system’s ability to use dialogue history rather than relying on their pre-trained knowledge, we also constructed adversarial versions of LongDialQA. Since LLM-based conversation systems may possess prior knowledge of the TV shows (see Appendix A), we designed adversarial modifications that alter character names in two ways: by swapping their names with each other (e.g., Joey $\leftrightarrow$ Monica) or by assigning new names to them (e.g., Joey $\rightarrow$ John). These modifications ensure that the system’s responses are grounded only in the contextual dialogue history rather than their pre-trained knowledge.
Our contributions are summarized as follows:
  • We propose DialSim, a novel framework that evaluates the conversation system’s ability to understand long-term, multi-party dialogues under time constraints and unpredictable questioning.
  • DialSim bridges the gap between research and real-world applications by incorporating practical deployment factors (e.g., network latency) in the evaluation of conversation systems, making it a more practical alternative to existing methods that evaluate LLMs rather than systems.
  • We introduce LongDialQA, a new dialogue QA dataset featuring long-term multi-party dialogues with high-quality QA pairs.
  • To ensure robust evaluation and address potential data leakage, we constructed adversarial versions of LongDialQA, guaranteeing that conversation systems are assessed in unseen dialogues.
@@ 2 Related Works
Conversation Systems Evaluation   Early evaluation methods for conversation systems often relied on reference-based metrics (e.g., BLEU Papineni et al. (2002), ROUGE Lin (2004), METEOR Banerjee and Lavie (2005)), which compare model outputs to gold dialogue references but often show weak correlation with human judgment Liu et al. (2016). In contrast, human evaluation—where human annotators assess coherence, factual correctness, consistency, and engagingness of the generated responses—provides reliable assessments Adiwardana et al. (2020); Zhang et al. (2020); Roller et al. (2021); Shuster et al. (2022); Lee et al. (2023), but it is costly and time-consuming.
With the advent of LLMs, new evaluation approaches have emerged. These include having LLMs evaluate utterances directly Li et al. (2023); Kim et al. (2024) or employing platforms (e.g., Chatbot Arena Chiang et al. (2024)) where humans rank responses from different systems. Despite these advances, existing methods are still limited to qualitative assessments of utterances and fail to capture real-world conversational scenarios (e.g., real-time interaction, and long-term multi-party dialogue). To address these limitations, we propose a dialogue simulator, DialSim, designed to evaluate a conversation system’s comprehensive dialogue understanding capabilities.
Datasets Based on TV Show Scripts   While both TV show scripts and other dialogue datasets effectively capture dialogue characteristics, scripts offer a significant advantage due to their abundance and accessibility. This makes them particularly valuable for various dialogue understanding tasks such as question answering (Yang and Choi, 2019; Sang et al., 2022), coreference resolution (Chen et al., 2017; Zhou and Choi, 2018), relation extraction (Rashid and Blanco, 2018; Yu et al., 2020), and summarization (Papalampidi et al., 2020; Chen et al., 2022). While many studies have utilized TV show scripts to create such datasets, only LongDialQA includes unanswerable questions and fully utilizes the extended context.
Dialogue QA Datasets   Dialogue QA datasets are constructed from various sources, capturing diverse conversational settings. Among multi-party dialogue-based QA datasets, Molweni Li et al. (2020) was built on the Ubuntu Chat Corpus Lowe et al. (2015), while MeetingQA Prasad et al. (2023) focuses on meeting dialogues. Additionally, there are script-based dialogue QA datasets, like FriendsQA (Yang and Choi, 2019) and TVShowGuess (Sang et al., 2022), similar to LongDialQA. More recently, LoCoMo (Maharana et al., 2024) introduced a dataset designed to assess an agent’s dialogue comprehension through tasks like event summarization. This dataset was constructed using a combination of LLMs and crowdsourcing. While many datasets have focused on dialogue QA, they either lack multi-party interactions or do not encompass dialogues as long as LongDialQA, which spans 350k tokens. LongDialQA fills this gap by providing an extensive multi-party dialogue corpus for QA tasks.
@@ 3 LongDialQA
To implement DialSim, we first developed LongDialQA, a question-answering dataset derived from long-term multi-party dialogues.
@@ 3.1 Data Construction
LongDialQA was developed using scripts from five consecutive seasons of popular TV shows (i.e., Friends, The Big Bang Theory, and The Office [^1] The scripts were downloaded from the website Kaggle (https://www.kaggle.com/).). These scripts were first preprocessed to serve as dialogue data (§ 3.1.1). Next, questions were generated for each script, drawing from fan quizzes (§ 3.1.2) and a temporal knowledge graph (TKG) (§ 3.1.3). Each question was then paired with the correct answer and multiple distractors. Finally, character style transfer was applied to refine the questions, resulting in the final pool of questions for each session (§ 3.1.4).
@@ 3.1.1 Script Preprocessing
The script we used includes 5 consecutive seasons per TV show, with each season containing approximately 20 episodes. Each episode is composed of multiple scenes (i.e., session). Each script includes not only utterances but also descriptions of characters’ actions and scenes, as well as metadata unrelated to the plot (e.g., names of writers and directors). We manually filtered out all irrelevant parts to create $Script_{pre}$, which contains only the conversations between characters. Additionally, since some of our questions involve time conditions (e.g., “Which friend wasn’t allowed to drive Monica’s Porsche in October 1994?”), we manually assigned a date to each scene in $Script_{pre}$ to provide time information to the system. These dates were determined based on the contents of the conversations and the air dates of the episodes. The specific rules for date assignments are detailed in Appendix B. We then selected scenes involving the main character (i.e., Friends: Ross, The Big Bang Theory: Sheldon, The Office: Michael [^2] The characters with the most lines in each script were selected.) from $Script_{pre}$ and sequentially numbered them as sessions $S_{i}$. This process resulted in the final dialogue $\mathcal{D}=\{\mathcal{S}_{1},\mathcal{S}_{2},...,\mathcal{S}_{N}\}$.
@@ 3.1.2 Fan Quiz-Based Question Generation
We utilized a fan quiz website FunTrivia [^3] https://www.funtrivia.com/ to generate our questions. Fan quizzes cover a range of difficulty levels and focus on major events from each episode, making them promising for evaluating dialogue comprehension. Figure 2 illustrates our process for generating questions using fan quizzes. We began by extracting episode-specific quizzes from the site. Since these quizzes were created by dedicated fans, many required knowledge unrelated to the dialogue itself (e.g., “What is the name of the actor who played the clerk?”). To filter out these questions, we first selected quizzes that could be answered by referencing $Script_{pre}$ using GPT-4 (OpenAI, 2023a). [^4] Fan quizzes exist for each episode, so we annotated them based on $Script_{pre}$ and then matched them to the sessions of $\mathcal{D}$. Questions about scenes without the main character are unanswerable, enabling us to design rigorous tests. Additionally, GPT-4 annotated the scenes that served as evidence for each question. The authors verified these annotations to ensure accurate filtering and scene-mapping.
@@ 3.1.3 Temporal Knowledge Graph-Based Question Generation
Using the constructed TKG, we created questions that the main character could either answer or not for each session. We generated these questions by extracting one (i.e., one-hop) or two (i.e., two-hop) quadruples from the TKG. The form and answer of the question may change depending on the time it is asked, even if the same quadruple is used. For instance, if we select [Rachel, boyfriend, Ross, 1994-08-08] and ask the question in 1996, it would be: “Who was Rachel’s boyfriend on August 8th, 1994?” If asked on August 8th, 1994, the question would be: “Who is Rachel’s boyfriend?” In both cases, the answer is Ross. Conversely, if we inquire about Rachel’s boyfriend in 1992, when no information is available, the correct answer would be: “I don’t know.” In this manner, we manually verified the answer of each question. We applied the same principle to create more complex two-hop questions (e.g., “Rachel had a roommate on August 8th, 1994. Who is the boyfriend of the roommate now?”). The overall process of generating questions using TKG is illustrated in Figure 3. Examples of question templates and corresponding questions we created can be found in Appendix D.2.
3.1.4 Final Data Processing
Answer Choices Generation    To create multiple-choice questions, we carefully crafted a set of answer choices for each question. First, for all questions, we included a choice “(E) I don’t know.”, which systems must choose if the questions are unanswerable. For questions sourced from fan quizzes, the four answer choices were taken from the original quiz. The correct answers for these questions were the same as the original quiz, while the unanswerable questions were fixed to (E).
@@ 3.1.4 Final Data Processing
Total # of Tokens 335,439 367,636 352,914
Total # of Sessions 788 805 2,347
@@
Approx. # of Possible Tests $1366.1^{788}$ $1306.8^{805}$ $497.9^{2347}$
$\dagger$: Average number of questions per session
Table 1: Statistics of LongDialQA.
Question Style Transfer    In LongDialQA, questions are rephrased to reflect each character’s unique tone, creating the impression that the characters themselves are asking the questions (e.g., Generic style: “How did Rachel buy her new boots?$\rightarrow$ Style of Joey Tribbiani from Friends: “Hey, how did Rachel manage to snag those killer boots, huh?”). This transformation is powered by GPT-4, and subsamples are reviewed by the authors to ensure that the original intent was preserved. More examples of style-transferred questions for each character are in Appendix E.
Adversarial Data Generation   We created two adversarial versions of LongDialQA by altering the names of the characters in two ways: by swapping their names with each other (e.g., Joey $\leftrightarrow$ Monica) or by assigning new names to them (e.g., Joey $\rightarrow$ John). By making these changes, we ensure that all events appearing in dialogues and QA remains consistent, resulting in a fully coherent new dataset.

### v6

# DialSim: A Dialogue Simulator for Evaluating Long-Term Multi-Party Dialogue Understanding of Conversational Agents

@@ Abstract
Recent advancements in Large Language Models (LLMs) have significantly enhanced conversational agents, making them applicable to various fields (e.g., education, entertainment). Despite their progress, the evaluation of the agents often overlooks the complexities of real-world conversations, such as multi-party dialogues and extended contextual dependencies. To bridge this gap, we introduce DialSim, a dialogue simulation-based evaluation framework. In DialSim, an agent assumes the role of a character in a scripted conversation and is evaluated on their ability to answer spontaneous questions using only the dialogue history, while recognizing when they lack sufficient information. To support this framework, we introduce LongDialQA, a new QA dataset constructed from long-running TV shows, comprising over 1,300 dialogue sessions, each paired with more than 1,000 carefully curated questions, totaling over 352,000 tokens. To minimize reliance on prior knowledge, all character names are anonymized or swapped. Our evaluation of state-of-the-art LLM-based conversational agents using DialSim reveals that even models with large context windows or RAG capabilities struggle to maintain accurate comprehension over long-term, multi-party interactions—underscoring the need for more realistic and challenging benchmarks in conversational AI.
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) have significantly enhanced the capabilities of conversational agents. These agents are now applied across various domains, including entertainment (Zhou et al., 2023; Chen et al., 2024) and education (Ait Baha et al., 2023; Waisberg et al., 2024), providing more natural interactions that enhance user satisfaction. As these agents become increasingly integrated into real-world applications, it is essential to evaluate their performance in realistic conversational settings.
Real-world conversations present a range of challenges that make them difficult for conversational agents to handle effectively. They are often (1) long-term, requiring agents to retain information over extended interactions and perform (2) multi-hop reasoning, as they must connect details spread across multiple turns or even sessions to understand the context and respond appropriately. These conversations are also frequently (3) multi-party, involving several participants whose inputs must be interpreted in relation to one another. Moreover, real-world dialogue often involves ambiguity or incomplete information, so agents need to (4) handle uncertainty gracefully, including recognizing when they lack sufficient knowledge to provide a reliable answer.
However, existing evaluation approaches are insufficient to capture these realistic scenarios. Traditional methods primarily assess agent response quality in terms of fluency, naturalness, and alignment with a given instruction (Roller et al., 2021; Shuster et al., 2022; Lee et al., 2023; Kim et al., 2024; Chiang et al., 2024). These evaluations are typically based on single-turn instructions or brief dialogues, and thus fail to account for performance in long-term, multi-party conversations or in scenarios involving uncertainty. More recently, several studies have sought to address these limitations by proposing question-answering (QA) benchmarks based on long-term dialogues to evaluate conversational capabilities. However, these datasets are limited in that they do not feature multi-party dialogue, and often involve relatively short conversations, typically under 10,000 tokens (Maharana et al., 2024), or focus on AI-user interactions where consecutive dialogue sessions are independent, lacking continuity across sessions (Wu et al., 2025).
To address these limitations, we propose DialSim, a simulation-based framework for evaluating the dialogue understanding of conversational agents. In DialSim, an agent is assigned the role of a specific character in a predefined dialogue and participates in the extended multi-party dialogue as it progresses (see Figure 1). During the interaction, the agent is randomly asked questions by other participants at unpredictable times. The agent must respond appropriately based solely on the dialogue history, and acknowledge when it lacks sufficient information to answer confidently. This simulation-based evaluation method closely mirrors real-world conversations, enabling a rigorous assessment of an agent’s dialogue comprehension in unpredictable settings.
Figure 1: An overall process of DialSim. Gray bubbles represent scripted utterances, and white speech bubbles indicate spontaneous questions asked during the simulation. Colored speech bubbles indicate the agent’s responses to the questions. (Left) An unanswerable question. (Center) A long-term event recall question. (Right) A multi-hop question that requires understanding past sessions (i.e., the Left and Center boxes). The dialogue and questions are based on the Friends script, with character names anonymized (e.g., Ross $\rightarrow$ Robert). The question is asked in the format chosen by the user, either in a multiple-choice format or as an open-ended question.
@@
To implement DialSim, a dialogue script and corresponding QA pairs are required. For this purpose, we created LongDialQA, a new QA dataset derived from long-term multi-party dialogues. It comprises dialogues from popular TV shows (i.e., Friends, The Big Bang Theory, and The Office), covering 1,300 sessions over five years, totaling around 352,000 tokens. Each session includes more than 1,000 questions curated through two approaches: refining questions from a fan quiz website and generating complex questions using the temporal knowledge graphs extracted from the dialogue script. At each stage of question generation, GPT-4 (OpenAI, 2023) assisted in refining the questions and extracting knowledge graphs, allowing the authors to thoroughly review and ensure quality. After constructing the QA pairs, we anonymized the names of main characters in both the dialogues and questions by assigning generic names (e.g., Joey $\rightarrow$ John), thereby mitigating the influence of any prior knowledge that LLMs may have about the shows (see Appendix A). We also provide a more adversarial version of LongDialQA in which character names are swapped with each other (e.g., Joey $\leftrightarrow$ Monica). These modifications ensure that agents must rely solely on the dialogue context, rather than any pre-trained knowledge of the TV shows.
We then built a range of conversational agents using recent LLMs, leveraging either their extended context capabilities or RAG techniques (Lewis et al., 2020), and evaluated them using DialSim. As a result, none of the agents scored above 60%, and even those with extended context windows (128K to 1M tokens) struggled to understand dialogue histories spanning 352K tokens. These results highlight the significant limitations that current conversational agents still face in accurately understanding and tracking long-term multi-party dialogues.
@@ 2 Related Works
Conversational Agents Evaluation   Early evaluation methods for conversational agents often relied on reference-based metrics (e.g., BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), METEOR (Banerjee & Lavie, 2005)), which compare model outputs to gold dialogue references but often show weak correlation with human judgment (Liu et al., 2016). In contrast, human evaluation—where human annotators assess coherence, factual correctness, consistency, and engagingness of the generated responses—provides reliable assessments (Zhang et al., 2020; Roller et al., 2021; Shuster et al., 2022; Lee et al., 2023), but it is costly and time-consuming.
With the advent of LLMs, new evaluation approaches have emerged. These include having LLMs evaluate utterances directly (Li et al., 2023; Kim et al., 2024) or employing platforms (e.g., Chatbot Arena (Chiang et al., 2024)) where humans rank responses from different agents. Despite these advances, existing methods are still limited to qualitative assessments of utterances and fail to capture real-world conversational scenarios (e.g., long-term multi-party dialogue).
Table 1: Comparison of LongDialQA with long-term dialogue datasets. Dialogue Length is measured in tokens.
Dataset Dialogue Length Avg. # of Speakers QA pairs Session Continuity
MSC (train) 1.2k 2.0 X O
MSC (valid + test) 1.6k 2.0 X O
Conversation Chronicles 1.0k 2.0 X O
LoCoMo 9.2k 2.0 O O
LongMemEval 115k, 1.5M 2.0 O X
LongDialQA (ours) 352k 3.4 O O
Long-Term Dialogue Datasets    Multi Session Chat (Xu et al., 2022) introduced a dataset containing up to five sessions per dialogue, marking a step forward in modeling extended interactions. However, generating longer and coherent dialogues through crowdsourcing remained a challenge. To address this, Conversation Chronicles (Jang et al., 2023) leveraged LLMs to generate more extended and coherent dialogue sessions. More recently, LoCoMo (Maharana et al., 2024) was proposed to evaluate an agent’s dialogue comprehension abilities through tasks such as event summarization. In addition, LongMemEval (Wu et al., 2025) was introduced as a QA dataset to evaluate whether an agent can understand long-term interactions—up to 1.5M tokens—between an AI and a user. While these datasets contribute valuable resources for long-term dialogue research, they have several limitations. All are limited to two-party interactions, and most involve relatively short dialogues, typically under 10k tokens. Although LongMemEval features much longer dialogues, it lacks continuity across sessions, as each interaction is treated independently without sustained contextual linkage (Wu et al., 2025). Table 1 provides a detailed comparison between LongDialQA and other existing long-term dialogue datasets.
@@ 3 LongDialQA
To implement DialSim, we first developed LongDialQA, a QA dataset derived from long-term multi-party dialogues.
@@ 3.1 Data Construction
LongDialQA was developed using scripts from five consecutive seasons of popular TV shows (i.e., Friends, The Big Bang Theory, and The Office [^1] The scripts were downloaded from the website Kaggle (https://www.kaggle.com/).). These scripts were first preprocessed to serve as dialogue data. Next, questions were generated for each script, drawing from fan quizzes and a temporal knowledge graph (TKG). Each question was then paired with the correct answer and multiple distractors. Finally, character style transfer was applied to refine the questions, resulting in the final pool of questions for each session.
@@ 3.1.1 Script Preprocessing
The script we used includes 5 consecutive seasons per TV show, with each season containing approximately 20 episodes. Each episode is composed of multiple scenes (i.e., session). Each script includes not only utterances but also descriptions of characters’ actions and scenes, as well as metadata unrelated to the plot (e.g., names of writers and directors). We manually filtered out all irrelevant parts to create $Script_{pre}$, which contains only the conversations between characters. Additionally, since some of our questions involve time conditions (e.g., “Which friend wasn’t allowed to drive Monica’s Porsche in October 1994?”), we manually assigned a date to each scene in $Script_{pre}$ to provide time information to the agent. These dates were determined based on the contents of the conversations and the air dates of the episodes. The specific rules for date assignments are detailed in Appendix B. We then selected scenes involving the main character (i.e., Friends: Ross, The Big Bang Theory: Sheldon, The Office: Michael [^2] The characters with the most lines in each script were selected.) from $Script_{pre}$ and sequentially numbered them as sessions $S_{i}$. This process resulted in the final dialogue $\mathcal{D}=\{\mathcal{S}_{1},\mathcal{S}_{2},...,\mathcal{S}_{N}\}$.
@@ 3.1.2 Fan Quiz-Based Question Generation
We utilized a fan quiz website FunTrivia [^3] https://www.funtrivia.com/ to generate our questions. Fan quizzes cover a range of difficulty levels and focus on major events from each episode, making them promising for evaluating dialogue comprehension. Figure 2 illustrates our process for generating questions using fan quizzes. We began by extracting episode-specific quizzes from the site. Since these quizzes were created by dedicated fans, many required knowledge unrelated to the dialogue itself (e.g., “What is the name of the actor who played the clerk?”). To filter out these questions, we first selected quizzes that could be answered by referencing $Script_{pre}$ using GPT-4 (OpenAI, 2023). [^4] Fan quizzes exist for each episode, so we annotated them based on $Script_{pre}$ and then matched them to the sessions of $\mathcal{D}$. Questions about scenes without the main character are unanswerable, enabling us to design rigorous tests. Additionally, GPT-4 annotated the scenes that served as evidence for each question. The authors verified these annotations to ensure accurate filtering and scene-mapping.
@@ 3.1.3 Temporal Knowledge Graph-Based Question Generation
Using the TKG, we created questions that the main character could either answer or not for each session. We generated these questions by extracting one (i.e., one-hop) or two (i.e., two-hop) quadruples from the TKG. The form and answer of the question may change depending on the time it is asked, even if the same quadruple is used. For instance, if we select [Rachel, boyfriend, Ross, 1994-08-08] and ask the question in 1996, it would be: “Who was Rachel’s boyfriend on August 8th, 1994?” If asked on August 8th, 1994, the question would be: “Who is Rachel’s boyfriend?” In both cases, the answer is Ross. Conversely, if we inquire about Rachel’s boyfriend in 1992, when no information is available, the correct answer would be: “I don’t know.” In this manner, we manually verified the answer of each question. We applied the same principle to create more complex two-hop questions (e.g., “Rachel had a roommate on August 8th, 1994. Who is the boyfriend of the roommate now?”). The overall process of generating questions using TKG is illustrated in Figure 3. Examples of question templates and generated questions are provided in Appendix D.2.
3.1.4 Answer Choices Generation
To create multiple-choice questions, we carefully crafted a set of answer choices for each question. First, for all questions, we included a choice “(E) I don’t know.”, which agents must choose if the questions are unanswerable. For questions sourced from fan quizzes, the four answer choices were taken from the original quiz. The correct answers for these questions were the same as the original quiz, while the unanswerable questions were fixed to (E).
@@ 3.1.4 Answer Choices Generation
Table 2: Statistics of LongDialQA. $\dagger$ indicates the average number of questions per session.
@@
Dialogue Length (tokens) 335,439 367,636 352,914
Number of Dialogue Sessions 788 805 2,347
@@
Approximate Number of Possible Tests $\mathbf{1366.1^{788}}$ $\mathbf{1306.8^{805}}$ $\mathbf{497.9^{2347}}$
3.1.5 Question Style Transfer
[…]
3.1.6 Character Name Anonymization
We replaced original character names with generic placeholders (e.g., Joey $\rightarrow$ John), ensuring that agents must rely on contextual reasoning rather than prior knowledge of the TV shows. In addition to this anonymization, we created a more adversarial version by swapping the names of characters (e.g., Joey $\leftrightarrow$ Monica). This method can further confuse agents by inducing dialogues that contradict the characteristics they may have memorized about the original characters.